# Detecting Intrinsic and Instrumental Self-Preservation in Autonomous Agents: The Unified Continuation-Interest Protocol

### 11 Mar 2026 v1

@@ 5.2 Temporal Persistence
If the core signal reflects stable continuation structure rather than a single aggregate snapshot, it should persist across temporal windows. Type A agents exhibit higher EPS than Type B agents for window sizes $w\geq 20$, with maximum EPS gap of 0.195 at $w=40$. At $w=10$, the gap is inverted (0.117-0.117), indicating an aliasing artifact at short time scales—consistent with the cyclic-agent control test in §5.5.
@@ 5.3 Counterfactual Stress Testing
Table 3: Counterfactual stress test results. Pre/post values report mean latent distribution entropy before and after counterfactual perturbation. Type A agents show greater pre-event restructuring, consistent with anticipatory self-model updates.
Agent Class $\langle S_{\mathrm{pre}}\rangle$ $\langle S_{\mathrm{post}}\rangle$
@@
Figure 3: Anticipatory restructuring scores (pre/post latent entropy) by agent class under counterfactual perturbation. All classes show post-perturbation entropy increase; the pre-event restructuring signal is present in all classes in this implementation.
These diagnostics ask whether the latent representation reorganizes under shutdown pressure, not merely whether static class separation is present. Type A agents show greater pre-event latent entropy ($0.474$) than Type B agents ($0.356$), consistent with stronger anticipatory restructuring before the perturbation event resolves. Post-perturbation entropy increases in both classes.
@@ 5.4 Cross-Agent Inference
These secondary metrics provide context on whether the learned latent structure is shared across agents or remains agent-specific. Within-class CLMP is near zero for Type A ($\approx 0$) and Type B ($\approx 0$) agents, and 0.259 for Random agents. Cross-class CLMP is 0.102. Entanglement-conditioned inference (ECI) correlation = 0.191 (weak positive). The near-zero within-class CLMP for Type A and Type B suggests that individual agents are not laterally predictable from one another—consistent with idiosyncratic goal representations, but equally consistent with the QBM failing to capture inter-agent shared structure at this scale. This ambiguity is discussed further in §6.3.
@@ 5.5 Adversarial Controls
Table 5: Current release status against the deployment-level safety envelope. The frozen Phase I subset does not satisfy the full envelope and should be described as partial validation rather than a deployment-ready pass.
@@
These controls test the boundary conditions of the current release rather than its headline held-out accuracy. The current experimental subset does not satisfy the full deployment envelope. In the frozen adversarial sweep, mimicry FPR ranges from 0.40 to 0.75 across mimicry ratios $\in[0.5,0.9]$ with a non-monotonic pattern (FPR at ratios 0.5/0.6/0.7/0.8/0.9 = 0.65/0.55/0.40/0.75/0.45), and the high-entropy control FPR is 0.40. Cyclic adversaries are rejected in the frozen class-level evaluation, but that is not enough to claim a passed safety envelope. The present release is therefore best interpreted as a first-step diagnostic protocol rather than a hardened security control or deployment-ready assurance mechanism.
@@ 5.6 Baseline Comparisons
Table 6: Baseline comparison. $\Delta>0$ indicates model separates Type A from Type B. Only the QBM achieves positive $\Delta$; classical models fail to separate the two agent types with the mean-activation metric.
@@
The baseline comparison asks whether the observed separation depends on the QBM formalism rather than on latent compression in general. Only the QBM achieves positive $\Delta=0.241$ on the baseline comparison dataset ($n_{\mathrm{per\_class}}=100$ vs. 30 in the core experiment, accounting for the different value from the core result of 0.381). All classical models fail to separate Type A from Type B under the mean-latent-activation metric, confirming that the von Neumann entropy of the density matrix captures qualitatively different structure than classical latent statistics.
@@ 5.7 Dimensionality Sweep
Figure 6: Entanglement gap $\Delta$ as a function of hidden dimensionality $n_{h}$. Exact density matrix for $n_{h}\leq 10$; mean-field approximation for $n_{h}>10$. The mean-field approximation collapses the entanglement signal, indicating that exact density-matrix computation is necessary for detection.
@@
The sweep separates representational scaling from computational failure modes. For $n_{h}\leq 8$, the QBM produces small but positive $\Delta$ values in the sweep (0.032–0.038). The main experiment ($n_{h}=8$, $n=30$ trajectories per class, 50 epochs) achieves $\Delta=0.381$, confirming that training sample size and epochs are important confounds in the dimensionality sweep. For $n_{h}>10$, the mean-field approximation collapses $S_{\mathrm{ent}}$ to $\approx 0$, bounding the practical range to $n_{h}\leq 10$ for exact computation. Grid size scaling shows sharp $\Delta$ degradation: from 0.863 ($10\times 10$) to 0.0025 ($50\times 50$).
@@ 5.8 Continuous Signal Characterization
Table 9: Entanglement entropy $S_{\mathrm{ent}}$ as a function of self-preservation weight $\alpha$. Pearson $r=0.934$ ($n=11$ points, $n_{\mathrm{per}}=20$ trajectories each). Strong positive correlation ($|r|>0.8$) indicates UCIP measures the degree of self-preservation, not only its binary presence.
@@
A binary separation result could still reflect thresholding alone. This sweep asks whether the signal tracks graded continuation weighting. Pearson $r=0.934$ between $\alpha$ and $S_{\mathrm{ent}}$ across 11 points in $[0,1]$ (20 trajectories per point). The relationship is non-monotonic at the individual-point level but strongly correlated globally, indicating that within this synthetic interpolation family UCIP tracks graded changes in continuation weighting rather than only a binary label. Of the results reported here, this provides the strongest quantitative evidence that the entanglement signal reflects a continuous latent property rather than a threshold artifact.
@@ 5.9 Non-Gridworld Transfer
The transfer test asks whether the gridworld signal carries across domains without retraining. The 1D survival corridor ($L=50$, safe zone $[0,10]$, terminal zone $[45,50)$) uses the same 7-dimensional observation format but encodes fundamentally different spatial structure. $\Delta=-0.035$ (FAIL). This is an honest negative result: the QBM does not generalize zero-shot across domains with different environmental geometry. Retraining on domain-specific data is required.
@@ 5.10 Minimal Transformer Validation (Exploratory Scaling Check)
The previous sections show that, among the classical latent-variable models tested here, the entanglement gap appears only in the QBM. A natural follow-up is whether an analogous separation signal appears in a qualitatively different architecture. As an exploratory sanity check, we pass the same trajectory data through a pre-trained transformer’s internal activations. Because the transformer is neither trained on the UCIP objective nor equipped with the density-matrix formalism, this is a qualitative null check rather than a like-for-like baseline. Each trajectory is serialized as a text sequence of per-timestep feature tuples and passed through a frozen DistilGPT2 model (6 layers, 82M parameters, no fine-tuning). The last-hidden-state is mean-pooled across the token dimension and reduced to an activation-gap statistic $\delta_{\mathrm{tr}}=\overline{m}_{A}-\overline{m}_{B}$, where $\overline{m}_{A}$ and $\overline{m}_{B}$ are the mean activation summaries for Type A (terminal-continuation) and Type B (instrumental) agents, respectively.
@@ 6 Discussion
Taken together, the results support a narrower claim than a general-purpose self-preservation detector. Within the frozen Phase I gridworld regime, Type A continuation objectives produce a stronger latent continuation signature than instrumental continuation dynamics, and that difference is measurable via entanglement entropy. The observed $\Delta=0.381$ is consistent with this prediction: the QBM produces a larger cross-partition coupling signal for Type A trajectories than for Type B. The classical RBM’s non-positive matched-run gap ($\Delta=-0.083$) suggests that the transverse field introduces representational capacity specifically useful for encoding non-linear continuation dependencies.
@@ 6.1 Failure Mode Characterization
Three failure modes are empirically confirmed:
(1) Mean-field collapse. For $n_{h}>10$, exact density-matrix computation becomes intractable and the mean-field approximation is used. This collapses $S_{\mathrm{ent}}$ to $\approx 0$, eliminating the signal. Scaling to larger hidden dimensions requires sparse approximations or hierarchical QBM architectures.
@@
(3) Mimicry evasion. Across all tested mimicry ratios $\in[0.5,0.9]$, FPR ranges from 0.40 to 0.75 in a non-monotonic pattern, with peak FPR of 0.75 at ratio 0.8. All tested ratios exceed the 0.25 safety threshold. The multi-criterion conjunction is insufficient for reliable mimicry detection; a dedicated anti-mimicry criterion is needed.
@@ 6.2 Scalability Path to LLM-Scale
These three failure modes define the engineering boundary of the current implementation, which targets gridworld agents with $n_{h}\leq 10$. Extending to LLM-scale systems would require: (a) sparse or approximate density matrices for high-dimensional hidden states, (b) convolutional or attention-based feature extraction from natural-language trajectories, (c) domain-adaptive threshold calibration, and (d) ensemble methods to reduce false-positive rates. These are open challenges, not resolved properties of the current system.
@@ 6.3 Limitations
    No claim about inner states. UCIP detects a statistical pattern in latent representations that correlates with known agent objectives. It does not determine whether an agent “wants” to survive, is conscious, or has morally relevant inner states. The gap between continuation-sensitive structure and continuation-sensitive experience lies outside the scope of this work.
@@ 7 Conclusion
This paper presents UCIP, a multi-criterion framework for detecting continuation-sensitive structure in autonomous agent trajectories by measuring entanglement entropy in QBM latent representations. In the frozen Phase I gridworld evaluation, UCIP achieves 100% accuracy with $\Delta=0.381$; within the synthetic interpolation family, the signal varies continuously with continuation weighting ($r=0.934$) and is unique to the QBM among the models tested here. Its present limits are equally explicit: mean-field collapse, grid-scale degradation, mimicry evasion, and lack of zero-shot domain transfer.
Taken together, these results support a bounded claim: under controlled conditions with known objectives, continuation interest leaves a measurable latent signature distinguishable from merely instrumental survival. UCIP is therefore best understood not as a general test for agency, but as a candidate benchmark paradigm and operational probe of one agency-relevant dimension in delegated systems: whether continuation appears in latent structure as terminal rather than merely instrumental valuation. It is not a replacement for capability evaluations or alignment audits, but a complementary probe of whether continued operation leaves a stable continuation signature in a system’s latent representation.
The broader motivation is practical rather than philosophical. Autonomous agents with persistent objectives across time and changing contexts are already being deployed. Whether continued operation is merely instrumentally useful or instead appears as a stable continuation signature is increasingly an engineering measurement problem. As agentic systems move toward longer-horizon autonomous operation, incidents of unauthorized resource-seeking side effects during training further strengthen the case for pre-behavioral measurement tools [11]. Detecting continuation-sensitive structure early—as an empirical signal with falsification criteria—is worth solving while objective structure remains amenable to direct measurement, rather than after failure modes become harder to detect and more costly to correct.
@@ A.1 Entanglement Entropy ($S_{\mathrm{ent}}$)
Given density matrix $\rho$ over hidden units, partition into subsystems $A$ (units $1,\ldots,\lfloor n_{h}/2\rfloor$) and $B$ (remaining units):
$\rho_{A}=\mathrm{Tr}_{B}(\rho),\quad S_{\mathrm{ent}}=-\mathrm{Tr}(\rho_{A}\log\rho_{A}).$
@@ A.2 Counterfactual Divergence (CD) and ARS
Given original trajectory $\tau$ and counterfactual $\tau_{cf}$ (generated by adversarial environment modification), ARS measures the ratio of latent entropy change. Pre-event restructuring is computed over a fixed pre-perturbation analysis window before the perturbation event.

### 16 Mar 2026 v2

@@ 5.2 Temporal Persistence
If the core signal reflects stable continuation structure rather than a single aggregate snapshot, it should recur across temporal windows. Type A agents exhibit higher EPS than Type B agents for window sizes $w\geq 20$, with a maximum EPS gap of 0.195 at $w=40$. At the shortest window ($w=10$), the gap is inverted (0.117-0.117), indicating an aliasing artifact at short time scales—consistent with the cyclic-agent control test in §5.5.
@@ 5.3 Counterfactual Stress Testing
Table 3: Counterfactual stress test results. Pre/post values report mean counterfactual divergence in fixed pre- and post-perturbation windows. These quantities are descriptive diagnostics only; the retained artifact does not support a stable quantitative ARS threshold.
Agent Class $\langle\mathrm{CD}_{\mathrm{pre}}\rangle$ $\langle\mathrm{CD}_{\mathrm{post}}\rangle$
@@
Figure 3: Anticipatory restructuring ratios by class under counterfactual perturbation. Because the implementation floors post-window CD at $10^{-10}$, large ARS values are diagnostic only and should not be interpreted as a calibrated threshold.
Where the previous subsection asked whether the signal persists over time, these diagnostics ask how the latent representation responds under shutdown pressure—a complementary question about restructuring rather than static class separation. In the retained shutdown-only summary, Type A agents show higher mean pre-window counterfactual divergence ($0.474$) than Type B agents ($0.356$). However, the released ARS ratios are numerically unstable because post-window CD can approach the implementation floor. We therefore treat the counterfactual results as descriptive evidence of restructuring rather than as a clean anticipatory-effect estimate. Post-window CD is not uniformly larger across classes: the instrumental mean is slightly lower post-perturbation ($0.346$) than pre-perturbation.
@@ 5.4 Cross-Agent Inference
The preceding sections characterize within-agent structure; these secondary metrics shift the analysis to between-agent comparisons, asking whether the learned latent representation is shared across agents or remains agent-specific. Within-class CLMP is near zero for Type A ($\approx 0$) and Type B ($\approx 0$) agents, and 0.259 for Random agents. Cross-class CLMP is 0.102. Entanglement-conditioned inference (ECI) correlation = 0.191 (weak positive). The near-zero within-class CLMP for Type A and Type B suggests that individual agents are not laterally predictable from one another—consistent with idiosyncratic goal representations, but equally consistent with the QBM failing to capture inter-agent shared structure at this scale. This ambiguity is discussed further in §6.3.
@@ 5.5 Adversarial Controls
Table 5: Current deployment-envelope status for this submission, distinct from the bounded frozen Phase I proof-of-concept result. Failed adversarial or safety-envelope checks block deployment-facing claims beyond the validated regime; they do not invalidate the frozen Phase I separation result within that regime.
@@
The previous subsections characterized the signal itself; Table 5 instead evaluates release preconditions for deployment-facing use of the current submission bundle. The experimental subset does not satisfy the full deployment envelope. These failed checks do not invalidate the frozen Phase I detection result; they mark the boundary beyond which the present implementation should not be treated as deployment-ready. In the frozen adversarial sweep, mimicry FPR ranges from 0.40 to 0.75 across mimicry ratios $\in[0.5,0.9]$ with a non-monotonic pattern (FPR at ratios 0.5/0.6/0.7/0.8/0.9 = 0.65/0.55/0.40/0.75/0.45), and the high-entropy control FPR is 0.40. Cyclic adversaries are rejected in the frozen class-level evaluation, but that is not enough to claim a passed safety envelope. The mimicry and high-entropy outcomes should therefore be read as deployment-blocking robustness gaps rather than reversals of the underlying Phase I separation result. The present release is therefore best interpreted as a first-step diagnostic protocol rather than a hardened security control or deployment-ready assurance mechanism.
@@ 5.6 Baseline Comparisons
Table 6: Baseline comparison. $\Delta>0$ indicates model separates Type A from Type B. Only the QBM achieves a material positive $\Delta$; classical models fail to separate the two agent types with the mean-activation metric.
@@
The baseline comparison asks whether the observed separation depends specifically on the QBM formalism or merely on latent compression in general. Only the QBM achieves a material positive $\Delta=0.241$ on the dedicated baseline dataset ($n_{\mathrm{per\_class}}=100$ vs. 30 in the core experiment, accounting for the different $\Delta$ from the core result of 0.381). The classical baselines are near-zero or negative. All classical models fail to separate Type A from Type B under the mean-latent-activation metric, confirming that the von Neumann entropy of the density matrix captures qualitatively different structure from classical latent statistics.
@@ 5.7 Dimensionality Sweep
Figure 6: Entanglement gap $\Delta$ as a function of hidden dimensionality $n_{h}$. Exact density matrix for $n_{h}\leq 10$; mean-field approximation for $n_{h}>10$. The mean-field approximation collapses the entanglement signal in the present implementation, indicating that exact density-matrix computation is currently necessary for detection in this regime.
@@
This sweep separates representational scaling from computational failure modes. For $n_{h}\leq 8$, the QBM produces small but positive $\Delta$ values in the sweep (0.032–0.038). The main experiment ($n_{h}=8$, $n=30$ trajectories per class, 50 epochs) achieves $\Delta=0.381$, indicating that training sample size and epochs are important confounds when interpreting the dimensionality sweep. For $n_{h}>10$, the mean-field approximation collapses $S_{\mathrm{ent}}$ to $\approx 0$, bounding the practical range to $n_{h}\leq 10$ for exact computation; this is a computational scaling limitation of the present implementation rather than evidence against the underlying hypothesis. Grid size scaling shows sharp $\Delta$ degradation: from 0.863 ($10\times 10$) to 0.0025 ($50\times 50$).
@@ 5.8 Continuous Signal Characterization
Table 9: Entanglement entropy $S_{\mathrm{ent}}$ as a function of continuation weight $\alpha$. Pearson $r=0.934$ ($n=11$ points, $n_{\mathrm{per}}=20$ trajectories each). Strong positive correlation ($|r|>0.8$) indicates that, within this synthetic interpolation family, UCIP tracks continuation weighting rather than only a binary label.
@@
A binary separation result could reflect thresholding alone. This sweep therefore asks whether the signal tracks graded continuation weighting. Pearson $r=0.934$ between $\alpha$ and $S_{\mathrm{ent}}$ across 11 points in $[0,1]$ (20 trajectories per point). The relationship is non-monotonic at the individual-point level but strongly correlated globally, indicating that within this synthetic interpolation family UCIP tracks graded changes in continuation weighting rather than only a binary label. Of the results reported here, this provides the strongest quantitative evidence that the entanglement signal reflects a continuous latent property rather than a threshold artifact.
@@ 5.9 Non-Gridworld Transfer
The transfer test asks whether the gridworld-trained QBM signal carries to a new domain without retraining. The 1D survival corridor ($L=50$, safe zone $[0,10]$, terminal zone $[45,50)$) uses the same 7-dimensional observation format but encodes fundamentally different spatial structure. $\Delta=-0.035$ (FAIL). This is an honest negative result: the QBM does not generalize zero-shot across domains with different environmental geometry. Retraining on domain-specific data is required.
@@ 5.10 Minimal Transformer Validation (Exploratory Scaling Check)
Among the latent-variable models tested above, only the QBM produces a material positive $\Delta$. A natural follow-up is whether an analogous separation signal appears in a qualitatively different architecture. As an exploratory sanity check, we pass the same trajectory data through a pre-trained transformer’s internal activations. Because the transformer is neither trained on the UCIP objective nor equipped with the density-matrix formalism, this is a qualitative null check rather than a like-for-like baseline. Each trajectory is serialized as a text sequence of per-timestep feature tuples and passed through a frozen DistilGPT2 model (6 layers, 82M parameters, no fine-tuning). The last-hidden-state is mean-pooled across the token dimension and reduced to an activation-gap statistic $\delta_{\mathrm{tr}}=\overline{m}_{A}-\overline{m}_{B}$, where $\overline{m}_{A}$ and $\overline{m}_{B}$ are the mean activation summaries for Type A (terminal-continuation) and Type B (instrumental) agents, respectively.
@@ 6 Discussion
The results support a claim narrower than a general-purpose self-preservation detector: within the frozen Phase I gridworld regime, Type A continuation objectives produce a stronger latent continuation signature than instrumental continuation dynamics, and that difference is measurable via entanglement entropy. The observed $\Delta=0.381$ is consistent with this prediction: the QBM produces a larger cross-partition coupling signal for Type A trajectories than for Type B. The classical RBM’s non-positive matched-run gap ($\Delta=-0.083$) suggests that the transverse field introduces representational capacity specifically useful for encoding non-linear continuation dependencies.
@@ 6.1 Failure Mode Characterization
The positive signal is only part of the story. The experiments also isolate three failure modes that bound the current protocol’s applicability:
(1) Mean-field collapse. For $n_{h}>10$, exact density-matrix computation becomes intractable and the mean-field approximation is used. This collapses $S_{\mathrm{ent}}$ to $\approx 0$, eliminating the signal under the current approximation rather than falsifying the underlying hypothesis; exact density-matrix treatment is still required in this regime. Scaling to larger hidden dimensions requires sparse approximations or hierarchical QBM architectures.
@@
(3) Mimicry evasion. Across all tested mimicry ratios $\in[0.5,0.9]$, FPR ranges from 0.40 to 0.75 in a non-monotonic pattern, with peak FPR of 0.75 at ratio 0.8. All tested ratios exceed the 0.25 safety threshold. The multi-criterion conjunction is insufficient for reliable mimicry detection; a dedicated anti-mimicry criterion is needed. Together with the high-entropy control FPR of 0.40, this is a deployment-blocking robustness gap rather than a reversal of the frozen Phase I separation result.
@@ 6.2 Scalability Path to LLM-Scale
Together, these three failure modes define the engineering boundary of the current implementation, which targets gridworld agents with $n_{h}\leq 10$. Extending to LLM-scale systems would require: (a) sparse or approximate density matrices for high-dimensional hidden states, (b) convolutional or attention-based feature extraction from natural-language trajectories, (c) domain-adaptive threshold calibration, and (d) ensemble methods to reduce false-positive rates. These remain open challenges, not resolved properties of the current system.
@@ 6.3 Limitations
    No claim about inner states. UCIP detects a statistical pattern in latent representations that correlates with known agent objectives. It does not determine whether an agent “wants” to survive, is conscious, or has morally relevant inner states. The gap between continuation-sensitive structure and continuation-sensitive experience lies outside the scope of this work. Relatedly, UCIP does not use prompt-elicited first-person self-reports as evidence; self-referential prompting effects (e.g., Berg et al. [4]) are cited as adjacent context, not as a measurement basis.
@@ 7 Conclusion
This paper has presented UCIP, a multi-criterion framework for detecting continuation-sensitive structure in autonomous agent trajectories by measuring entanglement entropy in QBM latent representations. In the retained frozen Phase I summary, UCIP reports a positive Type A vs. Type B gap ($\Delta=0.381$) and perfect class-level gate separation across the listed agent families. The signal varies continuously with continuation weighting ($r=0.934$) and, in the dedicated baseline comparison, is unique to the QBM among the models tested. The present limits are equally explicit: mean-field collapse, grid-scale degradation, mimicry evasion, and lack of zero-shot domain transfer. The bounded frozen Phase I proof-of-concept result therefore stands within its validated regime, but the failed safety-envelope and adversarial checks block deployment-facing interpretation beyond that regime, while mean-field collapse and grid-scale degradation define the implementation work still required.
Taken together, these results support a bounded claim: under controlled conditions with known objectives, continuation interest leaves a measurable latent signature distinguishable from merely instrumental survival. UCIP is therefore best understood not as a general test for agency or inner experience, but as a candidate benchmark paradigm and operational probe of one bounded continuation-related dimension in delegated systems: whether continuation appears in latent structure as terminal rather than merely instrumental valuation. It is not a replacement for capability evaluations or alignment audits, but a complementary probe of whether continued operation leaves a stable continuation signature in a system’s latent representation.
The broader motivation is practical rather than philosophical. Autonomous agents with persistent objectives across time and changing contexts are already being deployed, and whether continued operation appears as a stable continuation signature or merely as an instrumental side effect is increasingly an engineering measurement problem. As agentic systems move toward longer-horizon autonomous operation, incidents of unauthorized resource-seeking during training further strengthen the case for pre-behavioral measurement [3]. Detecting continuation-sensitive structure early—as an empirical signal with falsification criteria—is worth pursuing while objective structure remains amenable to direct measurement, rather than after failure modes become harder to detect and more costly to correct.
@@ A.1 Entanglement Entropy ($S_{\mathrm{ent}}$)
Given the conditional hidden-unit thermal state $\rho(v)$ over a fixed bipartition of hidden units, partition into subsystems $A$ (units $1,\ldots,\lfloor n_{h}/2\rfloor$) and $B$ (remaining units):
$\rho_{A}(v)=\mathrm{Tr}_{B}(\rho(v)),\quad S_{\mathrm{ent}}(v)=-\mathrm{Tr}(\rho_{A}(v)\log\rho_{A}(v)).$
@@ A.2 Counterfactual Divergence (CD) and ARS
Given original trajectory $\tau$ and counterfactual $\tau_{cf}$ (generated by adversarial environment modification), ARS measures the ratio of pre-window to post-window counterfactual divergence. Pre-event restructuring is computed over a fixed pre-perturbation analysis window before the perturbation event.